the average velocity at T1k and most propable velocity at T2k of co2 gas is9.0*104cm sec caluculate the values of T1and T2

Grade:9

1 Answers

Arun
6 years ago
Average velocity = √8RT/πM

and Most probable velocity = √2RT/M

Given –For CO2

Average velocity at T1 Most probable velocity at T2

= 9 * 104cm/sec = 9 * 104/100 m/sec.

= 9 * 102= m/sec.

∴ 9 * 102= √8 * 8.314 9 T1/3.14 *44 * 10-3 …..(A)

[Average velocity at T1K]

and 9 * 102= √2 * 8.314 * T2/44 * 10-3 …..(B)

[Most probable velocity at T2]

On solving, T1 = 1682.5 K, T2 = 2143.4 K